Find the length of segment RS if S between R and T, the length of RS is 1/3 the length of segment RT, RS=3x-3 and ST=2x+6.
Answer #1
RS + ST = RT
RT = 3RS
RS + ST = 3RS
3x-3 + 2x+6 = 3(3x-3)
5x + 3 = 9x - 9
4x = 12
x = 3

RS = 6
ST = 12
RT = 18
answered by: Adrian